Factor completely.
-3x^3 - 6x^2 + 189x

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

-3x^3 - 6x^2 + 189x
-3x(x^2+2x-63)=0
-3x(x^2+9x-7x-63)=0
-3x(x(x+9)-7(x+9))=0
-3x(x-7)(x+9)=0
(x-7)(x+9)=0/-3x
(x-7)(x+9)=0
x=7 or-9

Answer by ikleyn(53876)

.
Factor completely.
-3x^3 - 6x^2 + 189x
~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        The problem asks to factor completely.
        Instead, @mananth solved an equation, and got an incomplete solution.
        Thus, the requested assignment was not done in accordance with the request.

        I came to provide a correct solution.


  -3x^3 - 6x^2 + 189x = 

= -3x*(x^2 + 2x - 189) = 

= -3x*(x+9)*(x-7).     ANSWER

Solved.